# GPT-4V(ision) is a Generalist Web Agent, if Grounded

### v1

@@ Abstract
The recent development on large multimodal models (LMMs), especially GPT-4V(ision) and Gemini, has been quickly expanding the capability boundaries of multimodal models beyond traditional tasks like image captioning and visual question answering. In this work, we explore the potential of LMMs like GPT-4V as a generalist web agent that can follow natural language instructions to complete tasks on any given website. We propose SeeAct, a generalist web agent that harnesses the power of LMMs for integrated visual understanding and acting on the web. We evaluate on the recent Mind2Web benchmark. In addition to standard offline evaluation on cached websites, we enable a new online evaluation setting by developing a tool that allows running web agents on live websites. We show that GPT-4V presents a great potential for web agents—it can successfully complete % of the tasks on live websites if we manually ground its textual plans into actions on the websites. This substantially outperforms text-only LLMs like GPT-4 or smaller models (FLAN-T5 and BLIP-2) specifically fine-tuned for web agents. However, grounding still remains a major challenge. Existing LMM grounding strategies like set-of-mark prompting turns out not effective for web agents, and the best grounding strategy we develop in this paper leverages both the HTML text and visuals. Yet, there is still a substantial gap with oracle grounding, leaving ample room for further improvement.111Code and evaluation tool will be released at https://github.com/OSU-NLP-Group/SeeAct.
@@ 1 Introduction
Large multimodal models (LMMs; [22, 3, 26]), especially recent ones such as GPT-4V(ision) [31] and Gemini [4], have shown a remarkable capability on standard vision-and-language understanding and reasoning benchmarks [21, 15, 19, 37, 28, 56, 53]. While web content has been a primary source of training data, a largely overlooked part of the web is the websites themselves—every website is designed to be rendered visually for easy consumption by human users. This poses a new challenge and a new opportunity for LMMs. On the one hand, screenshots of rendered websites, which could contain thousands of elements with rich relations, are more complex than the images in most existing benchmarks, which are usually object- or scene-centric. On the other hand, if LMMs can accurately comprehend websites, it will open the door for numerous applications on the web.
@@
In this work, we aim to investigate the potential of LMMs as generalist web agents [12]. A generalist web agent, as defined in Mind2Web [12], is expected to follow natural language instructions and complete tasks on any given real-world website (e.g., Figure 1). The tasks can be fairly diverse and complex, with one task possibly taking + actions across multiple dynamically rendered webpages. Existing work [12, 27] primarily uses large language models (LLMs) such as GPT-4 [31] on the raw HTML input. However, HTML code is noisier than the rendered visuals and has a lower information density. For example, the screenshot in Figure 1 contains HTML elements that would require textual tokens with the GPT-2 Tokenizer, while requiring only visual tokens using GPT-4V’s visual tokenizer. Furthermore, HTML alone provides incomplete information and misses critical semantics from, e.g., embedded images.
@@
To this end, we propose SeeAct, a generalist web agent that harnesses the power of LMMs for integrated visual understanding and acting on the web. We will focus on GPT-4V, the most advanced LMM publicly available to date, and compare it with smaller LMMs such as BLIP-2 [22]. We find that GPT-4V exhibits a strong capability in visually understanding rendered webpages and generate the right plans in textual forms across a wide range of websites and tasks. However, grounding [6, 16], i.e., converting the textual plan into precise actions on the website, remains a major challenge. It involves selecting the right HTML element to interact with as well as the right operation (e.g., Click, Type, or Select). We propose multiple grounding methods, including superpositioning bounding boxes and index labels onto the image, similar to set-of-mark prompting [46] that has been shown effective on object- or scene-centric images. However, we find that on complex images with rich semantic and spatial relationships like webpage screenshots, severe hallucination is observed from GPT-4V. The most effective grounding strategy leverages the known correspondence between HTML element and their visual rendering, a unique property for websites compared to natural images.
@@
We evaluate SeeAct on the Mind2Web dataset [12] and compare it with text-only large language models (LLMs) like GPT-4 [31] as well as smaller models (FLAN-T5 [10] and BLIP-2 [22]) specifically fine-tuned for web agents. In addition to the standard offline evaluation setting on cached websites, we further establish an online evaluation setting by developing a new tool that allows for running web agents on live websites. The major findings from our exploration are summarized below:
@@
However, grounding is still a major challenge. The best grounding strategy still has a -% gap with oracle grounding. Among the various grounding strategies, the best one organically leverages both HTML text and visuals, substantially outperforming image annotation strategies [46] by up to %.
In-context learning with large models (both LMMs and LLMs) show better generalization to unseen websites, while supervised fine-tuning still has an edge on websites seen during training.
@@ 2 SeeAct
In this section, we first explain the problem formulation of web agent and then introduce our developed SeeAct, a generalist web agent based on GPT-4V.
Specifically, given a web-based task (e.g., “Rent a truck with the lowest rate” in the car rental website), we examine two essential capabilities of GPT-4V as a generalist web agent: (i) Action Generation to produce an action description at each step (e.g., “Move the cursor over the ‘Find Your Truck’ button and perform a click”) towards completing the task, and (ii) Element Grounding to identify an HTML element (e.g., “[button] Find Your Truck”) at the current step on the webpage.
@@ 2.1 Formulation
An action corresponds to a browser event provided by the website environment. Therefore, we formulate an action as a triplet of three necessary variables for a browser event . identifies the target webpage element to operate on, such as the "Find Your Truck" button in Figure 2.
represents the set of webpage elements within the environment . The operation is the action to be performed on the target element, with encompassing all possible operations in (e.g., Click, Type). The variable denotes the additional value needed for a certain operation (e.g., the date 12/10/2023 for a Type operation).
@@ 2.3 Action Grounding
To address this challenge, we explore three approaches using different types of information: Grounding via Element Attributes, Grounding via Textual Choices, and Grounding via Image Annotation, as depicted in Figure 2. The prompting details of action generation and grounding are included in Appendix A.
Grounding via Element Attributes. This approach involves prompting the model to generate as detailed attributes of the target element as possible, thereby providing more information to precisely match with the target HTML element.
Specifically, we prompt the model to not only describe the element , but also specify the target element’s type and the textual content in . For example, as illustrated in Figure 2, the model would generate element text as "Find Your Truck" and identify its type as a "BUTTON." Following this, a heuristic search is performed across the DOM elements, using the element text and type to locate matching elements. In cases where a single match is found, it is automatically selected. Otherwise, when multiple matches arise, the model is further prompted to select the final selection.
@@
Alternatively, we provide the model with textual representations of elements as choices to facilitate grounding, which has already been proven effective in MindAct [12]. Specifically, MindAct utilizes a ranker model to select top- candidate elements with a pretrained cross-encoder. Each candidate element is represented as a choice in a multi-choice question with its HTML text, as illustrated in Figure 2. After generating the action description , the model is further asked a multi-choice question to choose its intended element from the given multiple choices (including a ‘none’ option).
Grounding via Image Annotation. Textual representations alone are sometimes insufficient to distinguish similar or identical elements, as illustrated in Appendix E. Therefore, in this approach, we propose to overlay a bounding box for each candidate element selected by the ranker as well as an alphabet label on the bottom left of the bounding box222We use the Supervision library for image annotation: https://supervision.roboflow.com/.The model is expected to generate the label corresponding to the target element.
@@ 3.1 Dataset
We evaluate our methods on Mind2Web [12], a comprehensive dataset encompassing over complex web tasks with annotated actions. This dataset spans websites across low-level domains, categorized into high-level domains. It supports three primary operations: Click, Type, and Select, with Hover and Press Enter operations integrated into Click to avoid ambiguity.
@@
We align each HTML document in the dataset with its corresponding webpage screenshot image from the Mind2Web raw dump, which undergoes human verification to confirm correct rendering and element visibility. Given the constraints of the GPT-4V inference quota, our experiments focus on a random subset of 30 tasks from each test split. Detailed statistics of these subsets are presented in Table 1.
@@ 3.2 Methods
SeeAct. In grounding via image annotation and textual choices, we first employ the DeBERTa-base cross-encoder from MindAct [12] to rank the top 50 elements for better comparison with its text-only counterparts. Then, we cluster elements into groups of options for inference. In grounding via element attributes, no candidate element is provided. We access GPT-4V through ChatGPT WebUI, which only allows a zero-shot setting.
To compare with SeeAct, we also implement methods based on text-only LLMs and BLIP2 [22] following the two-stage strategy of MindAct [12]. Firstly, we employ the ranker above to pick the top 50 elements. Subsequently, the action generation problem is formulated as a multi-choice question answering problem, with the candidate elements as options, including a "None" option if the target element is absent. During inference, elements are clustered into groups of 5 elements, with iterative refinement, until a single choice is made or all options are discarded. We evaluate supervised fine-tuning (SFT) methods using FLAN-T5 [10] and BLIP2-T5 and in-context learning (ICL) methods using GPT-3.5, GPT-4.
FLAN-T5. We fine-tune FLAN-T5 using a left-to-right language modeling objective with the target sequence of ground-truth actions. The fine-tuned FLAN-T5 then serves as the backbone for inference, enabling action generation in the format defined in Equation 1.
BLIP2-T5. The BLIP-2 model combines a vision encoder and an LLM with a bridging component for modality connection. We jointly fine-tune the LLM and the bridge module on Mind2Web training data while keeping the vision encoder frozen. For the vision encoder, we leverage the ViT-L/14 pre-trained from CLIP [36] with an image resolution of . To ensure a fair comparison with the FLAN-T5-based text-only model, we choose FLAN-T5 as the language model and initialize it with the parameters fine-tuned on Mind2Web.
GPT-3.5 and GPT-4. We also conduct experiments with text-only LLMs, specifically GPT3.5-turbo-0613 and GPT4-turbo-1106-preview, using in-context learning in 3-shot settings. We use the same multiple-choice formulation and include three demonstration examples for in-context learning as specified in MindAct.
@@ 3.3 Offline Evaluation
We adopt the evaluation metrics utilized in Mind2Web. Element Accuracy (Ele. Acc) compares the predicted element with the ground-truth elements. Operation F1 (Op. F1) calculates the token-level F1 score for the predicted operation comprised of action and input value. Step Success Rate (Step SR) measures the success of each action step. A step is successful only if the selected element and the predicted operation are correct. We report macro averages across tasks for these step-wise metrics. Success Rate (SR) measures the success of an entire task. A task is regarded successful only if all steps have succeeded. This metric is stringent without allowing the model any space for exploration and error correction. Therefore, for offline evaluation, we focus on the first three metrics. However, we also conduct online evaluation on live websites to get a better evaluation on the whole task success rate, as detailed below.
@@ 3.4 Online Evaluation
We develop a new online evaluation tool using Playwright333https://playwright.dev/python/docs/api/class-playwright to evaluate web agents on live websites (instead of cached websites in offline evaluation). Our tool can convert the predicted action into a browser event. To adhere to ethical standards, our experiments are restricted to non-login tasks in compliance with user agreements, and we closely monitor agent activities during online evaluation to prevent any actions that have potentially harmful impact, like placing an order or modifying the user profile.
For a fair comparison between offline and online evaluations, we only re-write time-sensitive tasks to ensure they are still valid when the evaluation is conducted. For instance, we update the dates for flight-related tasks. Finally, we conduct the online evaluation on a subset of total tasks from thethree test splits and report the mean performance across these tasks.
@@
| FLAN-T5 | ||||||||||
| – Base | ||||||||||
| – Large | ||||||||||
| – XL | ||||||||||
| BLIP2-T5 | ||||||||||
| – Base | ||||||||||
| – Large | ||||||||||
| – XL | ||||||||||
@@
| GPT-3.5 | ||||||||||
| GPT-4 | ||||||||||
| SeeAct | ||||||||||
| – Attributes | ||||||||||
| – Annotation | ||||||||||
| – Choice | ||||||||||
| – Oracle | ||||||||||
@@ 4.1 Offline Evaluation Results
GPT-4V can be a Generalist Web Agent with Oracle Action Grounding. Given an effective action grounding method, GPT-4V has the potential to serve as a generalist web agent. Specifically, as described in subsection 2.3, we provide GPT-4V with an oracle action grounding method (SeeActOracle) through human annotation, the model achieves a step success rate of %, %, and % across three test splits, respectively. As shown in Table 2, this method substantially outperforms other models under all metrics across three test splits. Specifically, it achieves a % step success rate improvement over the second-best method in the Cross Task setting. The performance advantage is more pronounced under the Cross-Website and Cross-Domain settings, where it leads by 28.3% and 21.2% step success rates, demonstrating its generality compared with supervised fine-tuning. This observation is further corroborated with in the online evaluation (Table 3).
Element Grounding Method Comparison. However, there is a noticeable gap between oracle grounding and all the three proposed grounding methods. This demonstrates that grounding, especially element grounding, is a major bottleneck. Element grounding via textual choice (SeeActChoice) demonstrates the best performance under all metrics across all settings, comparable to supervised fine-tuning and showing a substantial improvement over text-only LLMs. Grounding via image annotation (SeeActAnnotation) offers an intuitive approach and shows promising results in recent work that focuses on object- or scene-centric images [46]. However, we find that on complex images with rich semantic and spacial relationships like webpage screenshots, severe hallucination is observed from GPT-4V. Specifically, it often fails to correctly map its generated element description (which is often correct according to oracle grounding) to the right bounding boxe and index label in the image, leading to a low element accuracy. This limitation primarily arises from GPT-4V’s weakness in understanding image details and relative spatial location, a topic that we will further delve into in subsection 4.3.
@@
LMMs vs. LLMs. The SeeActChoices model demonstrates a substantial performance advantage over the text-only GPT-4 under all three metrics across all three test splits. Specifically, it outperforms GPT-4 in step success rate of 8.9%, 14.7%, and 11.2% on three settings, respectively. Interestingly, fine-tuned BLIP2-T5 does not show a noticeable gain over FLAN-T5, despite having additional visual input. Several factors may contribute to this. First, the CLIP model used as the image encoder may not be sufficiently adept at image details, as explored by Shen et al. [38]. This limitation is particularly relevant for our web navigation task, which demands a high level of image detail comprehension. Second, BLIP2-T5 utilizes an off-the-shelf CLIP model that may not be optimal for webpage screenshots. Finally, although the screenshots in the test splits are error-free, some of the examples in the training set might contain issues such as rendering failures or inaccuracies in the timing of screenshot capture by annotators.
SFT vs. ICL. We compare SFT and ICL methods to offer insights for developing web agents in different scenarios. ICL (with SeeAct) demonstrate consistent and robust performance across three test splits. ICL is particularly advantageous in scenarios lacking annotations or requiring strong generalization capabilities for new domains and websites. As grounding methods improve towards oracle grounding, ICL is poised to show even stronger performance. On the other hand, SFT methods show better generalization across tasks on websites already seen during training. Considering the high cost of data annotation for web agents and the billions of websites on the internect, ICL offers a more compelling solution for generalist web agents. However, if one only needs to develop a strong web agent for a certain website, SFT is still a competitive solution.
4.2 Online Evaluation Results
In online evaluation, we pair a web agent with a human annotator, where the human was tasked to monitor agent actions that may change real-world states and determine whether each task was successfully completed. For comparative analysis, we include success rates from offline evaluation, denoted as Offline0 (allowing zero wrong action) and Offline1 (allowing one wrong action).

### v2

@@ Abstract
The recent development on large multimodal models (LMMs), especially GPT-4V(ision) and Gemini, has been quickly expanding the capability boundaries of multimodal models beyond traditional tasks like image captioning and visual question answering. In this work, we explore the potential of LMMs like GPT-4V as a generalist web agent that can follow natural language instructions to complete tasks on any given website. We propose SeeAct, a generalist web agent that harnesses the power of LMMs for integrated visual understanding and acting on the web. We evaluate on the recent Mind2Web benchmark. In addition to standard offline evaluation on cached websites, we enable a new online evaluation setting by developing a tool that allows running web agents on live websites. We show that GPT-4V presents a great potential for web agents—it can successfully complete % of the tasks on live websites if we manually ground its textual plans into actions on the websites. This substantially outperforms text-only LLMs like GPT-4 or smaller models (FLAN-T5 and BLIP-2) specifically fine-tuned for web agents. However, grounding still remains a major challenge. Existing LMM grounding strategies like set-of-mark prompting turns out to be not effective for web agents, and the best grounding strategy we develop in this paper leverages both the HTML structure and visuals. Yet, there is still a substantial gap with oracle grounding, leaving ample room for further improvement. All code, data, and evaluation tools are available at https://github.com/OSU-NLP-Group/SeeAct.
@@ 1 Introduction
Large multimodal models (LMMs; Li et al. (2023); Alayrac et al. (2022); Liu et al. (2023b)), especially recent ones such as GPT-4V(ision) (OpenAI, 2023) and Gemini (Anil et al., 2023), have shown a remarkable capability on standard vision-and-language understanding and reasoning benchmarks (Kazemzadeh et al., 2014; Goyal et al., 2016; Hendrycks et al., 2020; Saikh et al., 2022; Lu et al., 2022; Zhong et al., 2023; Yue et al., 2023). While web content has been a primary source of training data, a largely overlooked part of the web is the websites themselves—every website is designed to be rendered visually for easy consumption by human users. This poses a new challenge and a new opportunity for LMMs. On the one hand, screenshots of rendered websites, which could contain thousands of elements with rich relations, are more complex than the images in most existing benchmarks, which are usually object- or scene-centric. On the other hand, if LMMs can accurately comprehend websites, it will open the door for numerous applications on the web.
@@
In this work, we aim to investigate the potential of LMMs as generalist web agents (Deng et al., 2023). A generalist web agent, as defined in Mind2Web (Deng et al., 2023), is expected to follow natural language instructions and complete tasks on any given real-world website (e.g., Figure 1). The tasks can be fairly diverse and complex, with one task possibly taking + actions across multiple dynamically rendered webpages. Existing work (Deng et al., 2023; Liu et al., 2023d) primarily uses large language models (LLMs) such as GPT-4 (OpenAI, 2023) on the raw HTML input. However, HTML code is noisier than the rendered visuals and has a lower information density. For example, the screenshot in Figure 1 contains HTML elements that would require textual tokens with the GPT-2 Tokenizer, while requiring only visual tokens using GPT-4V’s visual tokenizer. Furthermore, HTML alone provides incomplete information and misses critical semantics from, e.g., embedded images.
@@
To this end, we propose SeeAct, a generalist web agent that harnesses the power of LMMs for integrated visual understanding and acting on the web. We will focus on GPT-4V, the most advanced LMM publicly available to date, and compare it with smaller LMMs such as BLIP-2 (Li et al., 2023) and LLaVA-1.5 (Liu et al., 2023a, c). We find that GPT-4V exhibits a strong capability in visually understanding rendered webpages and generate the right plans in textual forms across a wide range of websites and tasks. However, grounding (Chandu et al., 2021; Gu et al., 2023), i.e., converting the textual plan into precise actions on the website, remains a major challenge. It involves selecting the right HTML element to interact with as well as the right operation (e.g., Click, Type, or Select). We propose multiple grounding methods, including superpositioning bounding boxes and index labels onto the image, similar to set-of-mark prompting (Yang et al., 2023a) that has been shown effective on object- or scene-centric images. However, we find that on complex images with rich semantic and spatial relationships like webpage screenshots, severe hallucination is observed from GPT-4V. The most effective grounding strategy leverages the known correspondence between HTML element and their visual rendering, a unique property for websites compared to natural images.
@@
We evaluate SeeAct on the Mind2Web dataset (Deng et al., 2023) and compare it with text-only large language models (LLMs) like GPT-4 (OpenAI, 2023) as well as smaller models (FLAN-T5 (Chung et al., 2022), BLIP-2 (Li et al., 2023), LLaVA-1.5 (Liu et al., 2023a, c), and CogAgent (Hong et al., 2023). In addition to the standard offline evaluation setting on cached websites, we further establish a new online evaluation setting by developing a tool that allows for running web agents on live websites. The major findings from our exploration are summarized below:
@@
However, grounding is still a major challenge. The best grounding strategy still has a -% gap with oracle grounding. Among the various grounding strategies, the best one organically leverages both HTML text and visuals, substantially outperforming image annotation strategies (Yang et al., 2023a) by up to %.
In-context learning with large models (both LMMs and LLMs) shows better generalization to unseen websites, while supervised fine-tuning still has an edge on websites seen during training.
@@ 2 SeeAct
In this section, we first explain the problem formulation of web agents and then introduce SeeAct, a generalist web agent based on LMMs. Specifically, given a web-based task (e.g., “Rent a truck with the lowest rate” in the car rental website), we examine two essential capabilities of LMMs as a generalist web agent: (i) Action Generation to produce an action description at each step (e.g., “Move the cursor over the ‘Find Your Truck’ button and perform a click”) towards completing the task, and (ii) Element Grounding to identify an HTML element (e.g., “[button] Find Your Truck”) at the current step on the webpage.
@@ 2.1 Formulation
An action corresponds to a browser event provided by the website environment. Therefore, we formulate an action as a triplet of three necessary variables for a browser event . identifies the target webpage element to operate on, such as the "Find Your Truck" button in Figure 2. represents the set of webpage elements within the environment . The operation is the action to be performed on the target element, with encompassing all possible operations in (e.g., Click, Type). The variable denotes the additional value needed for a certain operation (e.g., the date 12/10/2023 for a Type operation).
@@ 2.3 Action Grounding
To address this challenge, we explore three approaches using different types of information: Grounding via Element Attributes, Grounding via Textual Choices, and Grounding via Image Annotation, as depicted in Figure 2. The prompting details of action generation and grounding are included in Appendix D.
Grounding via Element Attributes. This approach involves prompting the model to generate as detailed attributes of the target element as possible, thereby providing more information to precisely match with the target HTML element. Specifically, we prompt the model to not only describe the element , but also specify the target element’s type and the textual content in . For example, as illustrated in Figure 2, the model would generate element text as "Find Your Truck" and identify its type as a "BUTTON." Following this, a heuristic search is performed across the DOM elements, using the element text and type to locate matching elements. In cases where a single match is found, it is automatically selected. Otherwise, when multiple matches arise, the model is further prompted to select the final selection.
@@
Alternatively, we provide the model with textual representations of elements as choices to facilitate grounding, which has already been proven effective in MindAct (Deng et al., 2023). Specifically, MindAct utilizes a ranking model to select top- candidate elements with a pretrained cross-encoder. Each candidate element is represented as a choice in a multi-choice question with its HTML text, as illustrated in Figure 2. After generating the action description , the model is further asked a multi-choice question to choose its intended element from the given multiple choices (including a ‘none’ option).
Grounding via Image Annotation. Textual representations alone are sometimes insufficient to distinguish similar or identical elements, as illustrated in Appendix G. Therefore, in this approach, we propose to overlay a bounding box for each candidate element selected by the ranker as well as a label around the bounding box111We use the Supervision library for image annotation: https://supervision.roboflow.com/ with a label assignment method to avoid overlapping between markups. The model is expected to generate the label corresponding to the target element.
@@ 3.1 Dataset
We evaluate our methods on Mind2Web (Deng et al., 2023), a comprehensive dataset encompassing over complex web tasks with annotated actions. This dataset spans websites across low-level domains, categorized into high-level domains. It supports three primary operations: Click, Type, and Select, with Hover and Press Enter operations integrated into Click to avoid ambiguity.
@@
We align each HTML document in the dataset with its corresponding webpage screenshot image from the Mind2Web raw dump, which undergoes human verification to confirm element visibility and correct rendering for action prediction. This cleaned version of the dataset is called Multimodal Mind2Web222The dataset is released at https://huggingface.co/datasets/osunlp/Multimodal-Mind2Web., with the statistics in Table 1.
@@ 3.2 Methods
SeeAct. In grounding via image annotation and textual choices, we first employ the DeBERTa-base cross-encoder from MindAct (Deng et al., 2023) to rank the top 50 elements for better comparison with its text-only counterparts. Then, we cluster elements into groups of options for inference. In grounding via element attributes, no candidate element is provided. We experiment all three grounding methods with GPT-4V API, and use the best-performing grounding method for Gemini Pro Vision (Anil et al., 2023), and LLaVA-1.5 (Liu et al., 2023a, c).
MindAct. To compare with SeeAct, we also implement methods based on text-only LLMs and BLIP-2 (Li et al., 2023) following the two-stage strategy of MindAct (Deng et al., 2023). Firstly, we employ the ranker above to pick the top 50 elements. Subsequently, the action generation problem is formulated as a multi-choice question answering problem, with the candidate elements as options, including a "None" option if the target element is absent. During inference, elements are clustered into groups of 5 elements, with iterative refinement, until a single choice is made or all options are discarded. We evaluate supervised fine-tuning (SFT) methods using FLAN-T5 (Chung et al., 2022) and BLIP-2-T5 and in-context learning (ICL) methods using GPT-3.5 and GPT-4.
Pixel-Level Grounding. LMMs can generate target element coordinates in the image via training on datasets augmented with object coordinates, especially for open-sourced models (Hong et al., 2023; Cheng et al., 2024; You et al., 2023). We choose CogAgent (Hong et al., 2023) as a representative model for this experiment with the ICL method. Details of each method can be found in Appendix A.
@@ 3.3 Offline Evaluation
We adopt the evaluation metrics utilized in Mind2Web. Element Accuracy (Ele. Acc) compares the predicted element with the ground-truth elements. Operation F1 (Op. F1) calculates the token-level F1 score for the predicted operation comprised of action and input value. Step Success Rate (Step SR) measures the success of each action step. A step is successful only if the selected element and the predicted operation are correct. We report macro averages across tasks for these step-wise metrics. Success Rate (SR) measures the success of an entire task. A task is regarded successful only if all steps have succeeded. This metric is stringent without allowing the model any space for exploration and error correction. Therefore, for offline evaluation, we focus on the first three metrics. However, we also conduct online evaluation on live websites for better evaluation on the whole task success rate, as detailed below.
@@ 3.4 Online Evaluation
We develop a new online evaluation tool using Playwright333https://playwright.dev/ to evaluate web agents on live websites (instead of cached websites in offline evaluation). Our tool can efficiently tunnel multimodal inputs from the browser to the agent and convert the predicted action into a browser event for execution. To adhere to ethical standards, our experiments are restricted to non-login tasks in compliance with user agreements, and we closely monitor agent activities during online evaluation to prevent any actions that have potentially harmful impacts, like placing an order or modifying the user profile.
@@
| FLAN-T5-XL | ||||||||||
| BLIP-2-T5-XL | ||||||||||
@@
| GPT-3.5* | ||||||||||
| GPT-4* | ||||||||||
[…]
| – LLaVA-1.5 | ||||||||||
| – Gemini Pro Vision | ||||||||||
| – GPT-4V | ||||||||||
| – GPT-4V-Oracle* | ||||||||||
@@ 4.1 Offline Evaluation Results
GPT-4V can be a Generalist Web Agent with Oracle Action Grounding. Given an effective action grounding method, GPT-4V has the potential to serve as a generalist web agent. Specifically, as described in subsection 2.3, we provide GPT-4V with an oracle action grounding method (SeeActOracle) through human annotation, the model achieves a step success rate of %, %, and % across three test splits, respectively. As shown in Table 2, this method substantially outperforms other models under all metrics across three test splits. Specifically, it achieves a % step success rate improvement over the second-best method in the Cross-Task setting. The performance advantage is more pronounced under the Cross-Website and Cross-Domain settings, where it leads by 23.9% and 23.2% step success rates, demonstrating its generality compared with supervised fine-tuning. This observation is further corroborated within the online evaluation (Table 4).
Element Grounding Method Comparison. However, there is a noticeable gap between oracle grounding and all three proposed grounding methods, as shown in Table 3.
| Grounding | Cross-Task | Cross-Website | Cross-Domain |
| SeeActAttribute | |||
| SeeActAnnotation | |||
| SeeActChoice | |||
| SeeActOracle |
This demonstrates that grounding, especially element grounding, is a major bottleneck. Element grounding via textual choice (SeeActChoice) demonstrates the best performance under all metrics across all settings, comparable to supervised fine-tuning and showing a substantial improvement over text-only LLMs.
Grounding via image annotation (SeeActAnnotation) offers an intuitive approach and shows promising results in recent work that focuses on object- or scene-centric images (Yang et al., 2023a). However, we find that on complex images with rich semantic and spatial relationships like webpage screenshots, severe hallucination is observed from GPT-4V. Specifically, it often fails to correctly map its generated element description (which is often correct according to oracle grounding) to the right bounding box and index label in the image, leading to a low element accuracy. This limitation primarily arises from GPT-4V’s weakness in understanding image details and relative spatial location, a topic that we will further delve into in Appendix E. We leverage bottom-left number labels around bounding boxes as it is identified as the optimal markups in the ablation study in Appendix B.
@@
LMMs vs. LLMs. The SeeActChoice with GPT-4V demonstrates a substantial performance advantage over the text-only GPT-4 under all three metrics across all three test splits. Specifically, it outperforms GPT-4 in step success rate of %, %, and % on three settings, respectively. Interestingly, fine-tuned BLIP-2-T5 does not show a noticeable gain over FLAN-T5, despite having additional visual input. Several factors may contribute to this. First, the CLIP model used as the image encoder may not be sufficiently adept at image details, as explored by Shen et al. (2021). This limitation is particularly relevant for our web navigation task, which demands a high level of image detail comprehension. Second, BLIP-2-T5 utilizes an off-the-shelf CLIP model that may not be optimal for webpage screenshots. Finally, although the screenshots in the test splits are error-free, some of the examples in the training set might contain issues such as rendering failures or inaccuracies when annotators capture the screenshots.
@@
SFT vs. ICL. We compare SFT and ICL methods to offer insights for developing web agents in different scenarios. ICL (with SeeAct) demonstrates consistent and robust performance across three test splits. ICL is particularly advantageous in scenarios lacking annotations or requiring strong generalization capabilities for new domains and websites. As grounding methods improve towards oracle grounding, ICL is poised to show even stronger performance. On the other hand, SFT methods show better generalization across tasks on websites already seen during training. Considering the high cost of data annotation for web agents and the billions of websites on the Internet, ICL offers a more compelling solution for generalist web agents. However, if one only needs to develop a strong web agent for a certain website, SFT is still a competitive solution.
4.2 Online Evaluation Results
In online evaluation, we pair a web agent with a human annotator, where the human was tasked to monitor agent actions that may change real-world states and determine whether each task was successfully completed. For comparative analysis, we include success rates from offline evaluation, denoted as Offline0 (allowing zero wrong action) and Offline1 (allowing one wrong action). For a fair comparison between offline and online evaluations, we only re-write time-sensitive tasks to ensure they are still valid when the evaluation is conducted. For instance, we update the dates for flight-related tasks. Finally, we conduct the online evaluation on the same subset of total tasks from the three test splits. For tasks invalidated due to website change, we resample from the corresponding test split.